OpenStudy (anonymous):

A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies, how many did they sell each week? They sold boxes the first week, boxes the second week, and boxes the third week.

hero (hero):

Let: 1st Week = x 2nd Week = x + 5 3rd Week = 2(x+5) Then x + (x+5) + (2(x+5)) = 431

OpenStudy (anonymous):

Thanks

hero (hero):

You're welcome

hero (hero):

By the way, you still have to solve for x